OpenStudy (anonymous):

Why does obtuse traingle = a^2 > b^2 + c^2 and acute triangle = a^2 < b^2 +c ^2? What is a, b and c in each case? What exactly is < or > mean? Thanks! :)

OpenStudy (isaiah.feynman):

Any diagrams perhaps?

OpenStudy (isaiah.feynman):

< means "less than" > means "greater than"

OpenStudy (anonymous):

I saw someone answer this on another openstudy question. One moment, I'm trying to find it.

OpenStudy (campbell_st):

based on angles there are 3 types of triangles acute right obtuse so any triangle can be classified by applying pythagoras' theorem \[a^2 = b^2 +c^2 \] is a right triangle... the looking at what you posted you have the theorem fro obtuse and acute angled triangles.. hope it helps

OpenStudy (anonymous):

I understand right angled traingles but I don't understand obtuse or acute. I couldn't find the exact question, but let's use this previous question as a base: http://openstudy.com/study#/updates/50b6beb2e4b07527f06b1479

OpenStudy (anonymous):

7, 15 and 21 lengths - what triangle does it form?

OpenStudy (campbell_st):

ok so the longest side is the hypotenuse so \[21^2 = 441\] and \[7^2 + 15^2 = 274\] so \[21^2 > 7^2 + 15^2\] which means the triangle is obtuse

OpenStudy (anonymous):

Thank you Campbell, I understand now! Before I close my question, may I ask what is a acute triangle and what is obtuse?

OpenStudy (anonymous):

Thanks for helping too, Isaiah.Feynman :)